HUBBLE SPACE TELESCOPE DAILY REPORT #941 PERIOD COVERED: 0800 EST Daily Status Report as of 214/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C YSAFE (Safemode Recovery of the HST Science Instruments - FOS) The FOS Safemode recovery was nominal. No problems were reported. 1.2 Completed S/C XSAFE (Safemode Recovery of the HST Science Instruments - FOC) The FOC Safemode recovery was nominal. No problems were reported. 1.3 Completed S/C VSAFE (Safemode Recovery of the HST Science Instruments - HSP) The HSP Safemode recovery was nominal. No problems were reported. 1.4 Completed FOS/BL 4396 (QSO Absorption Systems Snapshot Survey - Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make one spectrographic observation of quasar 0253-0138 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The observation was executed as scheduled, and no problems were reported. 1.5 Completed Two Set of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled, and the observations were reported to look satisfactory. 1.6 Completed GHRS/2 3845 (Deciphering the UV Emission Lines in T Tauri Systems) The High Resolution Spectrograph was used to make six ultraviolet spectral observations of the T-Tauri class star BP-Tauri to study the width and shape of its emission lines which will put constraints on the type of mechanism producing these lines. One wavelength calibrations and six spectral observations were received and displayed. No problems were reported 1.7 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.8 Completed Two Sets of FOS/BL 4203 (Spatially Resolved Spectroscopy of Jupiter - Cycle 2/ April - May 1993) The Faint Object Spectrograph was used to make ultraviolet spectral observations of Jupiter's moon IO to make an "image spectrum" which was used to determine the position adjustment needed to center the spectrum of subsequent observations of Jupiter's red Spot and the North Pole Region. A corrective maneuver was successfully uplinked to the spacecraft. Io was reported to be well centered. Subsequent observations of Jupiter showed the Red Spot Near the limb of the Planet. The spectral observations looked fine. 1.9 Uplinked Updated Ephemris Tables, Per MOSES 2871 1.10 Completed Two Sets of PC 4809 (Cycle 2: WF/PC PSF Calibration Proposal) The Planetary Camera was used to make point spread function calibration image of standard star HD151406. The OSS monitored the observations and assessed the data quality. The last observations was reported to have been taken at an incorrect pointing location. 1.11 Completed GHRS/2 4601 (Jupiter's hydrogen Lyman Alpha Auroral/ Airglow Line Profile - Cycle 3 High) The High Resolution Spectrograph was used to make observations of the Hydrogen I emission line at 1216 A0 from Jupiter's atmosphere at latitudes of +40, +50, +55, +60 and +70 degrees along the central meridian. The observations were executed as scheduled, and no problems were reported. 1.12 Completed FOS/RD 5050 (Cycle 3 Internal Wavelenght Calibration) The Faint Object Spectrograph was used to make calibration spectra for every detector/ grating combination through the 0.3" circular and 0.1" paired apertures. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.13 Completed FOS/RD 4881 (Spect. Studies of Several High Z BAL-QSOs: Cycle 3 Medium) The Faint Object Spectrograph was used to make one observation of a broad absorption line quasar (Type BAL-QSO) known as Q1408+5642 as part of a survey of the ultraviolet spectra of these objects. The binary target acquisition failed because there was not enough counts in any of the read-out steps. The data was examined, and only noise was seen. The target apparently was not in the field of view. An HSTAR was filed. 1.14 Completed WFC 4849 (GTO High Latitude Parallel Program Cycle 3 High) The Wide Field Camera was used to make high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as planned, and no problems were reported. 1.15 Completed FOS/RD 4581 (Quasar Absorption Line Survey - Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make two observations of quasar Q1224-1116 to learn more about the origin and evolution of gas clouds along the line of sight of this target. Both observations were received and displayed. They had only noise, as was expected for this observations. No other anomalies were noted. 1.16 Completed FGS 4935 (Parallaxes of Hyades Cluster Members - Cycle 3 High) The Fine Guidance Sensors were used to make astrometric observations of star VA276 which is located in the Hyades cluster to obtain a definitive calibration of its population I zero-age main sequence members. The observation appeared to execute with no problems 1.17 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.18 Completed FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) The Faint Object Spectrograph was used to make two spectral observations of the planetary nebula LMC-SMP-87 which is located in the Large Magellanic Cloud. Both observations were received and displayed. The onboard target acquisition was successful, and the subsequent confirmation image showed that the target was well centered. 1.19 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. Both observations were received and displayed, and no problems were reported. 1.20 Completed FOS/RD/BL 4976 (UV Rotation Light Curves for Pluto, and Charon's UV Spectrum - Split) The Faint object Spectrograph was used to make a TALED exposure subsequent to other spectral observations of the planet Pluto and its moon Charon. The TALED observation was received in real time. It was reported to look okay. After a corrective SIP maneuver was successfully uplinked to the spacecraft, spectral observations were made of the planet Pluto. The observations were executed in fine lock. No problems were reported. 1.21 Completed Two Sets of GHRS/2 5039 (GHRS Dark Noise Monitoring 0 Cycle 3) The High Resolution Spectrograph was used to make twelve dark calibration images. The photoscan mode was used with an integration time of 10 seconds and with seven bins, for a read-out every 70 seconds. There was no carrousel motion. Observations were taken both inside and outside the small science aperture. All twelve observations were received and displayed. No anomalies were reported. 1.22 Completed FOS/RD 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: cycle 3 High) The Faint Object Spectrograph was used to make observations of quasar 2440+3539 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observations were executed as scheduled, and no problems were reported. 1.23 Completed FOS/BL 4600 (FUV Spectroscopy of Io's SO2 Atmosphere and Surface - Cycle 3 Medium) The Faint Object Spectrograph (Blue Side) was used to make spectral observations of the bright and dark regions on the surface of Jupiter's moon Io. A scan across Io was also made. The objective of this proposal was to further characterize the behavior fo SO2 in Io's atmosphere. The observations were not received in real-time. The analysis and offset links for this proposal were impacted by loss of the TDRS East satellite. The observations were received by the science tape recorder playback. After communications were re-established, several corrective maneuvers were uplinked to the spacecraft. Many of the following observations were reported to look satisfactory. One spectral observation was reported to have an noisy diode on channel 139. 1.24 Completed GHRS/2 5036 (GHRS 1ST Order Grating Spectral Lamp Test - Cycle 3) The High Resolution Spectrograph was used to maker eight spectral mini function calibration lamp tests of the first order gratings. All eight observations were received and displayed. in the OSS. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 19 Scheduled Reacquisitions : 22 Successful : 22 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 50 Successful : 50 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 4468 (Ultraviolet Spectropolarimetry of AG Car in its Current Outburst Cycle ) 3.2 Complete FOS/RD 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) 3.3 Complete FOS/BL 4976 (UV Rotation Light Curves for Pluto, and Charon's UV Spectrum - Split) 3.4 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-